# Privacy vs. Profit: The Impact of Google’s Manifest Version 3 (MV3) Update on Ad Blocker Effectiveness

### v1

@@ 7 Limitations and Ethics
Second, consistent with previous research [44], the automated browser-based experiment does not replicate all aspects of user interactions, such as visiting sub-pages of a website’s home page, which introduce different trackers than the website’s home page [50, 51, 52]. Future research could explore a broader range of user behaviors, including visiting sub-pages of a website’s home page, to mimic their real-world browsing behavior more closely.
Third, the study explores ad blocker effectiveness using a European IP address and default ad blocker settings on a select sample of websites focusing on display ads. Specifically, we examined four popular ad blockers, which does not represent the full ecosystem of ad blockers. Further, we focused on a select number of popular websites, which are likely skewed towards more popular ad vendors. We also primarily investigated header bidding ads served via Prebid framework, potentially missing other types of served ads (e.g., waterfall-style bidding). These choices provide a conservative estimate of ad blocker effectiveness, potentially missing the variability introduced by different geographic locations, ad blocker settings, and ad types, including video ads. Future work could examine these dimensions to uncover ad blocker effectiveness across diverse online environments and ad types.
Fourth, our study cannot exclude that any future changes to the Chrome extension ecosystem might hinder ad blocker effectiveness or that ad blocker providers get crowded out in the long term (e.g., through increased development and maintenance costs). Future examination of ad blocker effectiveness might therefore reveal nuanced results.
@@ 8 Conclusion
  • RQ1: Effectiveness Among MV3 and MV2 Ad Blocker Groups. The study revealed no statistically significant difference in the number of blocked ads or trackers between MV3 and MV2 groups of ad blockers. This result suggests that, contrary to concerns, the MV3 update did not reduce the ad-blocking or anti-tracking effectiveness of ad blockers.
@@
  • RQ2: Variability in Effectiveness Among Individual Ad Blockers. While the ad-blocking effectiveness remained consistent between MV3 and MV2 instances for most ad blockers, the Stands MV3 ad blocker blocked 6 more trackers than its MV2 counterpart, indicating a 30.5% improvement in anti-tracking effectiveness. This result suggests that the Stands ad blocker provider may have optimized their ad blocker to MV3 or that the MV3 update allows for enhanced anti-tracking effectiveness.
@@
  • RQ3: Enhanced Anti-Tracking Effectiveness with Multiple MV3 Ad Blockers. Contrary to the initial concerns of some ad blocker providers, using multiple MV3 ad blockers simultaneously did not reduce their ad-blocking effectiveness. Instead, it significantly enhances their anti-tracking effectiveness. Specifically, the combination of multiple MV3 ad blockers (Adblock Plus MV3, AdGuard MV3, Stands MV3, and uBlock MV3) blocks around 13 more trackers (61.8%) than Adblock Plus MV3 alone, 7 more trackers (25.4%) than AdGuard MV3 alone, and 7 more trackers (26.9%) than Stands MV3 alone. This improvement in anti-tracking effectiveness is driven by the inclusion of the uBlock MV3 ad blocker in the combination, and it is likely due to uBlock’s stronger anti-tracking default configurations after installation—as highlighted in related work [25, 26]. This finding indicates that concerns regarding the combined use of MV3 ad blockers were unfounded. Notably, this finding represents increased privacy and security for a privacy-aware user who uses multiple MV3 ad blockers simultaneously.
@@
In conclusion, this study contributes to understanding the impact of the MV3 update on ad blocker effectiveness. The findings indicate that MV3 ad blockers are as effective as their MV2 counterparts in blocking ads and trackers, and in some cases, may offer improved anti-tracking effectiveness. Results are consistent across different website samples, alternative effectiveness metrics, and over time. Moreover, cross-browser experiments yield comparable outcomes, and visual inspection confirms that MV3 ad blockers work effectively without significant ad flickering or loss of functionality. Still, MV3 ad blockers tend to offer a slightly less visually appealing browsing experience than their MV2 counterparts, mainly due to the increased visibility of cosmetic placeholders. These findings are reassuring for users who rely on ad blockers for a more private and ad-free browsing experience. The findings also highlight the successful adaptation of ad blocker providers to the MV3 update.
@@ A.1 Overview of Related Work on Ad Blocker Effectiveness
This study Number of blocked ads Number of blocked trackers $\checkmark$ $\checkmark$ All four ad blockers (Adblock Plus, AdGuard, Stands, uBlock) have equal ad-blocking effectiveness in their MV3 instances compared to their MV2 counterparts. Yet, the ad blocker Stands has improved anti-tracking effectiveness in its MV3 instance compared to its MV2 counterpart.
A.2 Variance Analysis for Browser-Based Experiment
In this section, we evaluate the consistency of our results across multiple runs of the browser-based experiment. Using variance analysis, we examine how much the results–specifically, the number of blocked ads and trackers–change when we repeat the runs of the browser-based experiment several times.
We use statistical measures of standard deviation (SD) and coefficient of variation (CV) to assess the consistency of the results. A low CV indicates that the results are consistent and do not vary much between runs. Table 4 presents the results of a variance analysis for ten separate measurement runs of our browser-based experiment involving a main sample of 114 websites.
Table 4: Variance analysis for the number of blocked ads and trackers across ten runs of browser-based experiment. CV means coefficient of variation. Multiplying the 9 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3, and MV3+) with the number of websites (114) yields the number of observations (N = 1,026).
@@ A.2 Variance Analysis for Browser-Based Experiment
Average SD 0.54 2.91
Average CV 0.21 0.25
No deviation 536 (52.24%) 170 (16.57%)
Max-min difference $\leq$ 1 727 (70.86%) 258 (25.15%)
N 1,026 1,026
Across these ten runs, the results are consistent, with coefficients of variation of 0.21 and 0.25 for the number of blocked ads and trackers, respectively. For blocked ads, in over half of the cases (52.24%), the number did not change between runs. In 70.86% of cases, the difference between the highest and lowest number of blocked ads was only one ad. The results for blocked trackers are similarly consistent, though absolute measures are higher due to a greater mean occurrence of trackers.
@@ B.1 Effectiveness of MV3 Ad Blockers Across Different Samples
To validate our individual ad blocker comparisons, we repeated our independent-sample t-test analyses on two separate stratified samples.
In our sample stratified by website employee count, we collected data from 191 websites and 9 browser instances, yielding 1,719 observations. In our sample stratified by Tranco popularity rank, we collected data from 185 websites and 9 browser instances, yielding 1,665 observations. For both stratified samples, we performed independent-sample t-tests comparing the MV3 to MV2 ad blockers (excluding the MV3+ ad blocker group) and then compared the MV3+ ad blocker group to the MV3 ad blockers.
@@
Figure 6: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group for sample stratified by website employee count. Panels A and C show independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (191) yields the number of observations (N = 1,719). Panels B and D show comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (191) yields the number of observations (N = 955). Error bars represent ±$\pm 1$ SEs.
Figure 7: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group for sample stratified by Tranco popularity rank. Panels A and C show independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (185) yields the number of observations (N = 1,665). Panels B and D show comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (185) yields the number of observations (N = 925). Error bars represent ±$\pm 1$ SEs.
For our sample stratified by website employee count, we observed no significant differences in the number of blocked ads between the MV3 and MV2 ad blockers, while for the number of blocked trackers, the Stands MV3 ad blocker outperformed its MV2 counterpart ($\Delta=7.03$, $p=0.03$). When comparing the MV3+ ad blocker group to the standalone MV3 ad blockers, we observed no significant differences in the number of blocked ads; however, for anti-tracking effectiveness (i.e., the number of blocked trackers) the MV3+ ad blocker group outperformed all standalone MV3 ad blockers except uBlock MV3.
@@ B.2 Alternative Measures for MV3 Ad Blockers Effectiveness
In this robustness test, we address limitations identified in Section 5.4 regarding our measures for ad blocker effectiveness.
We used the HTML file size (kb) of loaded websites under the usage of MV3 and MV2 ad blockers as an alternative, broader measure for ad-blocking and anti-tracking effectiveness of ad blockers. This measure follows the logic that successful blockage of unwanted ads and trackers, reduces the overall size of the loaded content.
@@
Table 5: HTML file size (kb) comparison for ad-blocking and anti-tracking effectiveness of ad blockers. Multiplying the number of websites (114) with the number of runs (10) for each browser instance (e.g., Adblock Plus MV2) yields the number of observations per browser instance (N = 1,140).
@@
Adblock Plus 1,140 884.77 623.90 886.34 622.47 -0.06 0.95
AdGuard 1,140 888.79 623.66 887.98 624.49 0.03 0.98
Stands 1,140 1113.29 642.82 904.72 638.39 7.77*** 0.00
uBlock 1,140 693.69 584.60 693.43 585.29 0.01 0.99
@@
Figure 8: Number of blocked third-party) domains between individual ad blockers and the MV3+ ad blocker group. Panel A a shows independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked domains. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (114) yields the number of observations (N = 912). Panel B shows comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked domains. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (114) yields the number of observations (N = 570). Error bars represent ±$\pm 1$ SEs.
In Panel A, we find no significant differences between the MV3 and MV2 ad blockers in the number of blocked domains for all ad blockers—except for Stands ($\Delta=10.06$, $p=0.02$). In Panel B, when comparing the MV3+ ad blocker group to the standalone MV3 ad blockers, the MV3+ ad blocker group blocks more domains than Adblock Plus MV3 ($\Delta=10.44,p<0.01$), AdGuard MV3 ($\Delta=3.78$, $p=0.08$), and Stands MV3 ($\Delta=3.90$, $p=0.07$), while uBlock MV3 shows no significant difference ($\Delta=0.02$, $p=0.99$). These results confirm a consistent pattern in anti-tracking effectiveness across both measures.

### v2

@@ 7 Limitations and Ethics
Second, consistent with previous research [52], the automated browser-based experiment does not replicate all aspects of user interactions, such as visiting sub-pages of a website’s homepage, which introduce different trackers than the website’s homepage [62, 63, 64]. We used the homepage as a uniform entry point for scale and comparability; most ad/analytics tags are installed site-wide, so we expect relative MV3 vs. MV2 differences to be similar across webpage types, even if some sub-pages may show higher numbers of ads or trackers. Future work could explore a broader range of behaviors, including visiting sub-pages.
Third, the study explores ad blocker effectiveness using a European IP address and default ad blocker settings on a select sample of websites, focusing on display ads. Specifically, we examined four popular ad blockers, which do not represent the full ecosystem of ad blockers. Further, we focused on a select number of websites, which are likely skewed towards more popular ad vendors. As a result, the web’s long tail is under-represented. Consequently, we cannot rule out that MV3’s 30,000-rule limit could disproportionately affect less popular websites that rely on rarely used rules. Fully testing the impact of this limit would require a larger set of less popular websites. We also primarily investigated header bidding ads served via Prebid framework, potentially missing other types of served ads (e.g., waterfall-style bidding). Because EU traffic often involves GDPR-driven website changes that reduce third-party activity irrespective of consent, baseline ad and tracker counts are lower. Thus, our absolute counts are conservative (lower-bound) estimates of ad blocker effectiveness. Future work could vary geography, ad blocker settings, and ad types (including video) to broaden external validity.
Fourth, our study cannot exclude that any future changes to the Chrome extension ecosystem might hinder ad blocker effectiveness, as feared by some ad blocker providers [15], or that ad blocker providers get crowded out in the long term (e.g., through increased development and maintenance costs). Future examination of ad blocker effectiveness might therefore reveal nuanced results.
@@ 8 Conclusion
  • RQ1: Effectiveness Among MV3 and MV2 Ad Blocker Groups. The study revealed no difference in the number of blocked ads between MV3 and MV2 groups of ad blockers. For trackers, the MV3 ad blocker group blocked about 1.8 more trackers on average per website than the MV2 group, indicating that the MV3 update did not reduce anti-tracking effectiveness.
@@
  • RQ2: Variability in Effectiveness Among Individual Ad Blockers. While the ad-blocking effectiveness remained consistent between MV3 and MV2 instances for most ad blockers, Adblock Plus MV3 blocked about 1.9 more trackers (21.5%) than its MV2 counterpart, and Stands MV3 blocked about 5.2 more trackers (45.9%) than its MV2 counterpart. This result suggests that these ad blocker providers may have optimized their ad blockers to MV3 or that MV3 allows for enhanced anti-tracking effectiveness.
@@
  • RQ3: Enhanced Anti-Tracking Effectiveness with Multiple MV3 Ad Blockers. Contrary to the initial concerns of some ad blocker providers, using multiple MV3 ad blockers simultaneously did not reduce their ad-blocking effectiveness. Instead, it significantly enhanced their anti-tracking effectiveness. Specifically, the combination of multiple MV3 ad blockers (Adblock Plus MV3, AdGuard MV3, Stands MV3, and uBlock MV3) blocked about 10.3 more trackers (95.3%) than Adblock Plus MV3 alone, 4.4 more (26.6%) than AdGuard MV3 alone, and 4.6 more (27.8%) than Stands MV3 alone, with no difference relative to uBlock MV3. The improvement in anti-tracking effectiveness is driven by the inclusion of the uBlock MV3 ad blocker in the combination, and it is likely due to uBlock’s stronger anti-tracking default configurations after installation—as highlighted in related work [26, 27]. This finding indicates that concerns regarding the combined use of MV3 ad blockers were unfounded. Notably, this finding represents increased privacy and security for a privacy-aware user who uses multiple MV3 ad blockers simultaneously.
In conclusion, this study contributes to understanding the impact of the MV3 update on ad blocker effectiveness. The empirical findings indicate no statistically significant reduction in ad-blocking or anti-tracking effectiveness for MV3 ad blockers compared to their MV2 counterparts. Some of our results even suggest slight improvements in blocking trackers for specific ad blockers following the MV3 update. These slight improvements might stem from filtering out unnecessary blocking rules (e.g. [14, 34]) or continued improvement of ad blockers by their providers.
@@
Our results hold across different website samples, alternative effectiveness metrics, and over time. Moreover, cross-browser experiments yield comparable outcomes, and visual inspection confirms that MV3 ad blockers work effectively without significant ad flickering or loss of functionality. Still, MV3 ad blockers tend to offer a slightly less visually appealing browsing experience than their MV2 counterparts, mainly due to the increased visibility of cosmetic placeholders. In total, these findings are reassuring for users who rely on ad blockers for a more private and ad-free browsing experience.
@@ A.1 Overview of Related Work on Ad Blocker Effectiveness
This study Number of blocked ads Number of blocked trackers $\checkmark$ $\checkmark$ All four ad blockers (Adblock Plus, AdGuard, Stands, uBlock) have equal ad-blocking effectiveness in their MV3 instances compared to their MV2 counterparts. Adblock Plus and Stands block more trackers under MV3 than under MV2, while AdGuard and uBlock do not; using multiple MV3 ad blockers improves anti-tracking relative to several single MV3 blockers.
A.2 Variance and Reliability Analysis for Browser-Based Experiment
In this section, we evaluate the consistency of our results across multiple runs of the browser-based experiment. Using variance analysis, we examine how much the results, specifically, the number of blocked ads and trackers, change when we repeat the runs of the browser-based experiment several times.
We use statistical measures of standard deviation (SD) and coefficient of variation (CV) to assess the consistency of the results. A low CV indicates that the results are consistent and do not vary much between runs. In addition, we compute the intraclass correlation coefficient (ICC), which measures how consistent the results are across runs by comparing the variation between different websites and ad blockers to the variation caused by repeating the browser-based experiment. Specifically, we report ICC(C,1)—a two-way, consistency, single-measure form of the intraclass correlation—together with 95% confidence intervals. A high ICC (>0.75) indicates high consistency across runs.
Table 4 presents the results of a variance and reliability analysis for five separate measurement runs of our browser-based experiment involving a main sample of 924 websites.
Table 4: Variance and reliability analysis for the number of blocked ads and trackers across five runs of browser-based experiment. CV means coefficient of variation. ICC means intraclass correlation coefficient and we report its values with 95% confidence intervals. Multiplying the 9 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3, and MV3+) with the number of websites (924) yields the number of observations (N = 8,316).
@@ A.2 Variance and Reliability Analysis for Browser-Based Experiment
Average SD 0.59 2.70
Average CV 0.67 0.20
No deviation 4,443 (53.43%) 1,581 (19.01%)
Max-min difference $\leq$ 1 5,612 (67.48%) 2,609 (31.37%)
ICC [95% CI] 0.77 [0.77, 0.78] 0.89 [0.88, 0.89]
N 8,316 8,316
Across these five runs, the results are consistent, with coefficients of variation of 0.67 and 0.20 for the number of blocked ads and trackers, respectively. For blocked ads, in over half of the cases (53.43%), the number did not change between runs. In 67.48% of cases, the difference between the highest and lowest number of blocked ads was only one ad. The results for blocked trackers are similarly consistent, though absolute measures are higher due to a greater mean occurrence of trackers. The ICC values confirm this impression: 0.77 for blocked ads and 0.89 for blocked trackers, indicating good and excellent consistency, respectively.
@@ B.1 Effectiveness of MV3 Ad Blockers Across Different Samples
To validate our individual ad blocker comparisons, we repeated our independent-sample t-test analyses on two separate stratified samples. First, we stratified by website employee count. We excluded missing values and treated each observed employee count (0, 1, 10, 100, 1,000, 100,00) as a stratum. Within each such stratum, we sampled without replacement up to a fixed number of websites. Second, we stratified by Tranco popularity rank. We recoded the "Outside Top 1m" rank to 1,000,000, removed missing values, and partitioned the resulting distribution into ten equal-frequency bins (deciles). We then sampled a fixed number of websites per decile without replacement.
In our sample stratified by website employee count, we collected data from 191 websites and 9 browser instances on October 29, 2024, yielding 1,719 observations. In our sample stratified by Tranco popularity rank, we collected data from 185 websites and 9 browser instances on October 29, 2024, yielding 1,665 observations. For both stratified samples, we performed independent-sample t-tests comparing the MV3 to MV2 ad blockers (excluding the MV3+ ad blocker group) and then compared the MV3+ ad blocker group to the MV3 ad blockers.
@@
Figure 6: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group for sample stratified by website employee count. Panels A and C show independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (191) yields the number of observations (N = 1,719). Panels B and D show comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (191) yields the number of observations (N = 955). Values shown are from a single measurement run (no averaging); error bars represent $\pm 1$ SE across websites.
Figure 7: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group for sample stratified by Tranco popularity rank. Panels A and C show independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (185) yields the number of observations (N = 1,665). Panels B and D show comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (185) yields the number of observations (N = 925). Values shown are from a single measurement run (no averaging); error bars represent $\pm 1$ SE across websites.
For our sample stratified by website employee count, we observed no significant differences in the number of blocked ads between the MV3 and MV2 ad blockers, while for the number of blocked trackers, the Stands MV3 ad blocker outperformed its MV2 counterpart ($\Delta=7.03$, $p<0.001$). When comparing the MV3+ ad blocker group to the standalone MV3 ad blockers, we observed no significant differences in the number of blocked ads; however, for anti-tracking effectiveness (i.e., the number of blocked trackers) the MV3+ ad blocker group outperformed all standalone MV3 ad blockers except uBlock MV3.
@@ B.2 Alternative Measures for MV3 Ad Blockers Effectiveness
In this robustness test, we address limitations identified in Section 5.4 regarding our measures for ad blocker effectiveness in our main sample of 924 websites.
We used the HTML file size (kb) of loaded websites under the usage of MV3 and MV2 ad blockers as an alternative, broader measure for ad-blocking and anti-tracking effectiveness of ad blockers. This measure follows the logic that successful blockage of unwanted ads and trackers reduces the overall size of the loaded content.
@@
Table 5: HTML file size (kb) comparison for ad-blocking and anti-tracking effectiveness of ad blockers. Multiplying the number of websites (924) by the two versions compared for each blocker (MV2 and MV3) yields the number of observations across both versions (1,848). Values shown are per-website averages across five runs.
@@
Adblock Plus 924 485.95 501.39 485.23 510.30 -0.03 0.98
AdGuard 924 447.08 504.37 447.78 503.28 0.03 0.98
Stands 924 671.67 506.36 450.28 505.74 -9.40 0.00
uBlock 924 423.12 470.34 420.65 466.43 -0.11 0.91
@@
Figure 8: Number of blocked third-party domains between individual ad blockers and the MV3+ ad blocker group. Panel A shows independent t-test comparisons between the mean values of the MV3 and MV2 ad blockers for the number of blocked domains. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 8 browser instances (Adblock Plus MV2, Adblock Plus MV3, AdGuard MV2, AdGuard MV3, Stands MV2, Stands MV3, uBlock MV2, uBlock MV3) with the number of websites (924) yields the number of observations (N = 7,392). Panel B shows comparisons between the MV3+ ad blocker group and the individual MV3 ad blockers for the number of blocked domains. Multiplying the 5 browser instances (Adblock Plus MV3, AdGuard MV3, Stands MV3, uBlock MV3, and MV3+ ad blocker group) with the number of websites (924) yields the number of observations (N = 4,620). Values shown are per-website averages across five runs; error bars represent $\pm 1$ SE across websites.
In Panel A, we find no significant differences between the MV3 and MV2 ad blockers in the number of blocked domains for AdGuard and uBlock; however, Adblock Plus ($\Delta=2.70$, $p<0.001$) and Stands ($\Delta=9.07$, $p<0.001$) block more domains under MV3 than under MV2. In Panel B, when comparing the MV3+ ad blocker group to the standalone MV3 ad blockers, the MV3+ ad blocker group blocks more domains than Adblock Plus MV3 ($\Delta=15.17,p<0.001$), AdGuard MV3 ($\Delta=7.19$, $p<0.001$), and Stands MV3 ($\Delta=6.94$, $p<0.001$), while uBlock MV3 shows no significant difference. These results confirm a consistent pattern in anti-tracking effectiveness across both measures.